### v1

# Deep Support Vectors

Abstract
While the success of deep learning is commonly attributed to its theoretical equivalence with Support Vector Machines (SVM), the practical implications of this relationship have not been thoroughly explored. This paper pioneers an exploration in this domain, specifically focusing on the identification of Deep Support Vectors (DSVs) within deep learning models. We introduce the concept of DeepKKT conditions, an adaptation of the traditional Karush-Kuhn-Tucker (KKT) conditions tailored for deep learning. Through empirical investigations, we illustrate that DSVs exhibit similarities to support vectors in SVM, offering a tangible method to interpret the decision-making criteria of models. Additionally, our findings demonstrate that models can be effectively reconstructed using DSVs, resembling the process in SVM. The code will be available.
Machine Learning, ICML
1 Introduction
Deep learning has led revolutionary advancements in modern technology, catalyzing paradigm shifts across numerous fields. However, despite these successes, a fundamental understanding of how deep learning models excel in pattern generalization and the criteria they use for decision making remains uncertain. This research endeavors to demystify the black box of deep learning, offering deeper insights into how these models process and interpret data.
In explaining the performance of deep learning, one of the most useful attempts is associating it with support vector machines (SVM) (Cortes & Vapnik, 1995; Hearst et al., 1998) because SVM theoretically provides a well-tight bound on its performance. Consequently, support vectors, which encode the decision boundary, offer valuable insights into how the model classifies based on certain criteria. Moreover, the number of support vectors is much smaller than the size of the trained dataset, making it parameter-efficient. This idea aligns well with deep learning models that converge to 100% training accuracy in overparameterized settings (Zhang et al., 2021). It has been proven that feedforward networks, including convolutional networks, can be equivalent to SVM (Lyu & Li, 2020; Soudry et al., 2018; Chen et al., 2021) and there is an active effort to bring the SVM architecture itself into the realm of deep learning models (Wang et al., 2019; Tang, 2013).
Hence, it can be logically inferred that a deep learning model, akin to an SVM, should demonstrate comparable fundamental traits. Specifically, the model is expected to have a corresponding set of support vectors that define its decision boundary. However, attempts to extract support vectors from generally trained deep learning models have not yet been made.
In this paper, we focus on the crucial issue of the practical relationship between support vectors and deep learning. Based on the theoretical equivalence (Soudry et al., 2018), our research represents the first attempt to connect traditional machine learning, epitomized by SVM, with typical deep learning models. In doing so, we define ‘Deep Support Vectors’ (DSVs) as the samples constituting decision boundary in the input space and their desirable characteristics analogous to their counterparts in linear SVMs. Reinterpreting the traditional SVM’s KKT (Karush-Kuhn-Tucker) condition (Bertsekas, 1999) for deep networks, we introduce ’DeepKKT’, a revised KKT condition tailored for deep learning, along with the concept of Deep Support Vectors. DeepKKT smoothly blends the principles of classical machine learning with the deep learning realm, making it compatible with the deep learning feature of handling high-dimensional data that reside on a low-dimensional manifold.
Figure 1: Comparison of synthesized images (first row) created using the DeepKKT condition initiated from noise, and selected images (second row) from the CIFAR-10 training dataset. The selected images were chosen based on $\lambda$ values, i.e., the ten images have the highest $\lambda$ in each class. Both synthesized and selected images demonstrate similarity at the pixel level sharing common features.
In our experiments with DSVs, we utilize DeepKKT in two ways. Firstly, we select DSVs from the training dataset using DeepKKT as a scoring function measuring the closeness of a sample to the decision boundary. Secondly, we synthesize DSVs from pure noise, employing DeepKKT as a loss function in model inversion settings (See Fig. 1).
In this study, we performed experiments to elucidate the attributes of synthesized DSVs. Our findings reveal that DSVs consistently meet all conditions analogous to support vectors in SVMs. Additionally, we demonstrated that DSVs contain significantly denser information than other data samples in the training dataset. Furthermore, we demonstrated that modifying the predictions of the model by editing the information gathered from DSVs is possible. This illustrates that DSVs can play a role in elucidating the decision criteria of the model. Also, we demonstrated that DSVs can actually serve as the same role as support vectors, meaning than we can retrain the model up to a level using only a single image per category. i.e., we used DSVs in dataset distillation (Cazenavette et al., 2022; Wang et al., 2018; Liu et al., 2022). Our performance is comparable to others for preexisting benchmarks . Also, our setting is far more reasonable than others because they demand substantial computational resources as they normally use gradient matching between the distilled dataset and the entire training dataset. This requires heavy computational load, which is paradoxical as their goal is to reduce computation burden while maintaining privacy.
In contrast to existing studies that solely provide input-dependent explanations for decision-making, DSVs are the first method capable of elucidating the decision boundary’s criterion solely through the trained model, thereby unraveling the black-box nature of deep learning.
2 Related Works
2.1 SVM in Deep Learning
Numerous studies have endeavored to establish a connection between deep learning and conventional SVMs. Recently, Soudry et al. (2018) demonstrated that, in an overparameterized setting with linearly separable data, the learning dynamics of a model possessing a logistic tail are equivalent to those of a support vector machine, with the model’s normalized weights converging towards a finite value. Following this, Lyu & Li (2020) extended this equivalence to feedforward networks. Other studies (Wang et al., 2019; Tang, 2013) directly integrated SVM into their architecture to capitalize on its beneficial properties. Despite these advancements, there is still a lack of research that directly explores support vectors through this theoretical lens of equivalence. In this study, we address this theoretical equivalence of support vectors by introducing the DeepKKT condition, a KKT condition tailored for deep learning, thus applying this equivalence in a practical context.
2.2 Model Inversion Through the Lens of Maximum Margian
There is a line of research utilizing the stationarity condition, a part of the KKT condition, for model inversion. Haim et al. (2022) firstly exploited the KKT condition for model generation, adopting SVM-like architectures. They conducted experiments with binary classification, a 2-layer MLP, and full-batch gradient descent. Buzaglo et al. (2023) extended these experiments to multi-label classification by adapting their existing architecture to a multi-class SVM structure. To ensure the generated samples lie on the data manifold, they were initialized with the dataset’s mean, implying the adoption of some prior knowledge of the data. Similarly, Yu & Wang (2023) generated images through the stationarity condition, also adopting the mean-initialization and conducting experiments on the MNIST (Deng, 2012) and CelebA (Liu et al., 2015) datasets.
Contrary to previous studies, our research extensively employs the KKT conditions, not just focusing on the stationarity condition. While previous works adopted SVM-like structures to directly utilize the stationarity condition, we derived DSVs from universal deep learning settings including SGD (Stochastic Gradient Descent), mini-batch processing, multi-class classification, and data augmentation, indicating that DSVs can be obtained in any trained model settings, not just within SVM-like architectures. Furthermore, their limited number aligns with the core principle of support vectors. These distinctions underscore the unique contributions of our work. This was possible because we reinterpreted the KKT condition used in machine learning to fit in deep learning settings.
2.3 Dataset Distillation
Dataset distillation (Cazenavette et al., 2022; Wang et al., 2018; Liu et al., 2022) fundamentally aims to reduce the size of the original dataset while maintaining model performance. The achievement also involves addressing privacy concerns and alleviating communication issues between server nodes. Although this approach does not explicitly clarify the link with SVM, distilled datasets share key features with support vectors in that they enable model reconstruction with minimal data. The problem is typically addressed under the following conditions: 1) Access to the entire dataset for gradient matching, 2) Possession of snapshots from all stages of the model’s training phase, and 3) Availability of multiple models with different initializations. These conditions present greater complexity and constraints relative to our DSV extraction, posing practical implementation challenges in real-world situations. In contrast, our algorithm does not require all three of these conditions and operates based just on a single trained model.
3 Preliminaries
3.1 Notation
In SVM formulation, $\tilde{w}(:=[w;b])$ represents the concatenated weight vector $w$ and bias $b$. Each data instance, expanded to include the bias term, is denoted by $\tilde{x}_{i}(:=[x_{i};1])$, while the corresponding binary label is represented by $y_{i}\in\{\pm 1\}$. The Lagrange multipliers are denoted by $\alpha_{i}$’s.
Transitioning to the context of deep learning, we denote the parameter vector of a neural network model by $\theta$, which, upon optimization, yields $\theta^{*}$ as the set of learned weights. The mapping function $\Phi(x_{i};\theta)$ represents the transformation of input data into a $C$-dimensional logit in a $C$-class classification problem in a manner dictated by the parameters $\theta$, i.e., $\Phi(x_{i};\theta)=[\Phi_{1}(x_{i};\theta),\cdots,\Phi_{C}(x_{i};\theta)]^{T}\in\mathbb{R}^{C}$. we define score as the logit of a target index. For example, $\Phi_{y_{i}}(x_{i};\theta)$ is the value of the logit for the target class. If the score is the largest among logits, i.e., $\operatorname*{argmax}_{c}\ \Phi_{c}(x_{i};\theta)=y_{i}$, then it correctly classifies the sample. The Lagrange multipliers adapted to the optimization in deep learning are represented by $\lambda_{i}$’s. $(x,y)$ denotes a pair of input and output and $\mathcal{I}$ is the index set.
This section presents the specific conditions that DSVs (Deep Support Vectors) must satisfy and discusses how to achieve that condition within gradient descent framework.
3.2 Support Vector Machines
The fundamental concept of Support Vector Machines (SVMs) is to find the optimal hyperplane that classifies the given data. This hyperplane is defined by the closest data points to itself known as support vectors, and the distance between the support vectors and the hyperplane is termed the margin. The hyperplane must classify the classes correctly while maximizing the margin. This leads to the following conditions that an SVM must satisfy:
Primal feasibility: $\forall i,\ y_{i}\tilde{w}^{T}\tilde{x}_{i}\geq 1,$ (1)
Dual feasibility: $\forall i,\ \alpha_{i}\geq 0,$
[…]
Stationarity: $\tilde{w}=\sum_{i=1}^{n}\alpha_{i}y_{i}\tilde{x}_{i}.$
The primal and dual conditions ensure that these critical values correctly classify the data while being outside the margin. The complementary slackness condition mandates that support vectors lie on the decision boundary. The stationarity condition makes sure that the derivative of the Lagrangian be zero.
The final condition, stationarity, offers profound insights into Support Vector Machines (SVM). It underscores the fact that support vectors are integral to encoding the decision boundary $\tilde{w}$. Consequently, identifying the decision hyperplane in SVMs is tantamount to pinpointing the corresponding support vectors. This implies that with a trained model at our disposal, we can reconstruct the SVM in two distinct ways:
1. Support Vector Selection: From the trained model, we can extract support vectors that inherently encode the decision hyperplane.
2. Support Vector Synthesis: Alternatively, it is feasible to generate or synthesize support vectors, even in the absence of a training set, which can effectively represent the decision hyperplane by generating samples that satisfy $|\tilde{w}^{T}\tilde{x}|=1$.
4 Deep Support Vector
This section presents the specific conditions that DSVs (Deep Support Vectors) must satisfy and discusses how to get an optimization loss with low computation cost that meets these conditions.
Algorithm 1 Support Vector Refinement for Deep Learning Model
  Input: Pretrained model $\Phi(\cdot;\theta)$, loss function $L$, augmentation function set $A$, number of DSV candidate $N$, number of class $C$, hyperparameter $\beta$, $\gamma$, $\alpha_{\lambda}$, $\alpha_{x}$
  Ensure: Freeze model $\Phi(\cdot;\theta)$
  Initialize $N\times C$ number of support vector candidates
  for $c=1$ to $C$ do
     Sample $N$ number of ($x^{c}_{i}$,$\lambda^{c}_{i}$) for label $c$
  end for
  Define $X=\{x^{c}_{i}\mid c\in[C],i\in[N]\}$
  Define $\Lambda=\{\lambda^{c}_{i}\mid c\in[C],i\in[N]\}$
  repeat
     $L_{\text{primal}}(X)=(\sum_{i=1}^{N}\sum_{c=1}^{C}L(\Phi(x^{c}_{i};\theta),c))/NC$
     $L_{\text{stat}}(X)=||\theta+\sum_{i=1}^{N}\sum_{c=1}^{C}\lambda^{c}_{i}\nabla_{\theta}L(\Phi(x^{c}_{i};\theta),c)||^{2}_{2}$
     $L_{\text{kkt}}(X)=L_{\text{primal}}(X)+\beta\times L_{\text{stat}}(X)$
     Sample $f_{A}\in A$
     Define $X^{\prime}=\{f_{A}(x^{c}_{i})\mid x^{c}_{i}\in X\}$
     ${L_{total}(X)}={L_{kkt}(X)}+\gamma\times{L_{kkt}(X^{\prime})}$
     Update $x_{i}^{c}\leftarrow x_{i}^{c}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{x}}\nabla_{x_{i}^{c}}L_{total}(X)$,  $\forall\ i\ \text{and}\ c$
     Update $\lambda_{i}^{c}\leftarrow\lambda_{i}^{c}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{\lambda}}\nabla_{\lambda_{i}^{c}}L_{total}(X)$,  $\forall\ i\ \text{and}\ c$
     Remove $x^{c}_{i}$ from $X$ if corresponding $\lambda^{c}_{i}<0$
  until $X$ converges
  Return: Set of DSV : $X$
4.1 DeepKKT
SVM’s Relationship with Hinge loss
We start our discussion by focusing on the hinge loss, a continuous surrogate loss for the primal feasibility, and its gradient:
$\begin{split}&\text{Hinge Loss:}\ L_{h}(x_{i},y_{i};\tilde{w})=\frac{1}{n}\sum_{i=1}^{n}\max(0,1-y_{i}(\tilde{w}^{T}\tilde{x}_{i})),\\ &\quad\nabla_{\tilde{w}}L_{h}=\frac{1}{n}\sum_{i=1}^{n}\begin{cases}0,&\text{if }y_{i}(\tilde{w}^{T}\tilde{x}_{i})>1\\ -y_{i}\tilde{x}_{i},&\text{otherwise.}\end{cases}\end{split}$ (2)
With Eq. (LABEL:eq:hingegrad), the stationarity condition in Eq. (1) becomes
$w^{*}:=-\sum_{i=1}^{n}\alpha_{i}\nabla_{w}L(x_{i},y_{i};w^{*}),\quad\text{s.t.}\ \alpha_{i}>0.$ (3)
Generalization of conventional KKT conditions
In this paper, we adapt the KKT conditions presented in Eq. (1) to deep learning. In doing so, the two main hurdles of a deep network different from a linear binary SVM are 1) the nonlinearity of $\Phi(x_{i};\theta)$ taking the role of $\tilde{w}^{T}\tilde{x}_{i}$ and 2) multi-class nature of a deep learning model.
Considering the role of the primary condition is to correctly classify $x_{i}$ into $y_{i}$, we can enforce the score $\Phi_{y_{i}}(x_{i};\theta)$ for the correct class $y_{i}$ to take the maximum value among all the logits with some margin $\epsilon$, i.e.,
$\Phi_{y_{i}}(x_{i};\theta^{*})-\max_{c\neq y_{i},c\in[C]}\Phi_{c}(x_{i};\theta^{*})\geq\epsilon,$ (4)
Because this condition is discontinuous, we can relax it with a surrogate function $-L(\Phi(x_{i};\theta^{*}),y_{i})$, which is the negative loss function to maximize. Note that if we take the cross-entropy loss for $L$, it becomes
$-L_{ce}=\Phi_{y_{i}}(x_{i};\theta^{*})-\log\sum_{c=1}^{C}\exp(\Phi_{c}(x_{i};\theta^{*})),$ (5)
which takes a similar form as Eq. (4) and the negative loss can be maximized to meet the condition.
Now that we found the analogy between $y_{i}\tilde{w}^{T}\tilde{x}_{i}$ and $-L(\Phi(x_{i};\theta^{*}),y_{i})$, $y_{i}\tilde{x}_{i}(=\nabla_{\tilde{w}}(y_{i}\tilde{w}^{T}\tilde{x}_{i}))$ corresponds to $-\nabla_{\theta}^{*}L(\Phi(x_{i};\theta^{*}),y_{i})$. Thus, the stationary condition $\tilde{w}=\sum_{i=1}^{n}\alpha_{i}y_{i}\tilde{x}_{i}$ in SVMs can be translated into that of deep networks such that
$\theta^{*}=-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}).$ (6)
This condition is also a generalized formulation of Eq. (3), where we substitute the linear model $\tilde{w}^{T}\tilde{x}$ with a nonlinear model $\Phi(x;\theta)$, and the binary classification hinge loss $L_{h}$ with multi-class classification loss $L$. Furthermore, the stationarity condition reflects the dynamics of overparameterized deep learning models. We provide an analogy in Appendix A.
However, these conditions are not enough for deep learning. As mentioned before, we are interested in synthesizing support vectors. Compared to the problems dealt with in the classical SVMs, the input dimensions of deep learning problems are typically much higher. In this case, the data are likely to lie along a low-dimensional latent manifold $\mathcal{M}$ inside the high-dimensional space. To ensure a generated sample be a plausible DSV, it not only should satisfy the generalized KKT condition but also should lie in an appropriate data manifold, i.e., $x\in\mathcal{M}$.
Finally, we can rewrite the new DeepKKT condition as follows:
Primal feasibility: $\forall\,i\in\mathcal{I},\quad\operatorname*{argmax}_{c}\;\Phi_{c}(x_{i};\theta^{*})=y_{i}$
Dual feasibility: $\forall\,i\in\mathcal{I},\quad\lambda_{i}\geq 0,$
Stationarity: $\theta^{*}=-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}\mathcal{L}(\Phi(x_{i};\theta^{*}),y_{i}),$
Manifold: $\forall\,i\in\mathcal{I},\quad x_{i}\in\mathcal{M}.$
4.2 Deep Support Vectors, From Dust to Glory
In Sec, 4.1, we explored the KKT conditions in the context of deep learning and how these conditions can be used to formulate a loss function. It is important to note that our goal is not to construct the model $\Phi$, but rather to generate support vectors of an already-trained model $\Phi$ with its parameter $\theta^{*}$.
Our paper addresses the issue of reconstructing support vectors from trained deep learning models. Essentially, this involves sampling or synthesizing support vectors. By replacing the optimization variable $\theta$ with $x$, we shift our focus to utilizing the DeepKKT conditions for generating or evaluating input $x$ rather than $\theta$. This adjustment necessitates a consideration of the data’s inherent characteristics, specifically its multiclass nature and the tendency of the data to reside on a lower-dimensional manifold compared to its defined dimension.
Primal Feasibility
Firstly, the primal feasibility condition which is equivalent to Eq. (4) questions the fundamental performance of the model. That is, the trained model $\theta^{*}$ must be able to correctly discriminate its support vectors. As presented in Sec. 4.1, instead of Eq. (4), we use a surrogate function for the loss:
$L_{i}=\begin{cases}0&\text{if }\operatorname*{argmax}_{c}\Phi_{c}(x_{i};\theta^{*})=y_{i},\\ L(\Phi(x_{i};\theta^{*}),y_{i})&\text{otherwise}.\end{cases}$
This is designed to match the primal condition by mimicking the Hinge loss. When each DSV is correctly classified, no loss is incurred. Otherwise, we adjust the DSVs to align them with the correct target, effectively optimizing the support vectors. This approach also implicitly enforces the complementary slackness condition, as $L_{i}$ does not promote increased confidence in the incorrect classification. Consequently, the final loss, which refines the support vectors, is computed as follows:
$L_{\text{primal}}=\frac{1}{n}\sum_{i=1}^{n}L_{i}.$ (7)
Here, $L$ can be any loss function and we have employed the cross-entropy loss in our experiments.
Stationarity
Secondly, the stationarity condition can be used directly as a loss function. Since we are extracting DSVs from the trained model $\Phi(\cdot;\theta^{*})$, we construct this loss as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{\text{stat}}=D(\theta^{*},-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}))}$ (8)
For the distance measure $D$, any metric can be used; we have chosen to use the $l_{1}$ distance to suppress the effect of outliers. It is crucial to remember that our objective is to find DSVs and the optimization is done for the primal and dual variables $x_{i}$ and $\lambda_{i}$ and not for the parameter $\theta$. For this, we require one forward pass and two backward passes; one for $\nabla_{\theta}L$ and the other for $\nabla_{x_{i}}D$. The overall computational cost is quite low, as we optimize only a small number of samples.
Moreover, as shown in Algorithm 1, we satisfy the dual condition by ensuring the Lagrange multipliers $\lambda_{i}$’s are greater than zero and disqualify any $x_{i}$’s from being a support vector candidate if during optimization $\lambda_{i}$ becomes less than zero. The condition that is not directly satisfied is the complementary slackness. To directly fulfill the functional boundary for support vectors as specified in Sec. 3.2, we would need to be able to calculate the distance between functions, which is not only abstract but also requires a second-order computation cost. Therefore, we adopted a relaxed version of the KKT conditions that excludes this requirement. Furthermore, as demonstrated in Sec. 5.2, we have shown that DSVs implicitly satisfy the complementary slackness condition. This implies that DSVs meet every condition introduced in conventional SVM.
Figure 2: Conceptual overview of our algorithm. Initially sampled from noise, the DSVs evolve to intersect with the KKT solution subspace and the data manifold.
Manifold Condition: Reflecting High-Dimensional Dynamics of Deep Learning
In addition to satisfying the conventional KKT conditions, we have adopted additional criteria for determining DSVs. As much as determining the data distribution in a high-dimensional space, giving additional constraints will allow DSVs to approach the desired data manifold. The DeepKKT conditions form an DSV solution subspace as inferred in Fig. 2, however this does not guarantee that it is on the data manifold that the data is trained on. Among the DSVs that satisfy the DeepKKT conditions, we must select the ‘good’ DSVs closest to the manifold. Since the model is trained on the data manifold, it follows that the DSVs would also be formed within this manifold. By selecting DSVs that are in the intersection of the solution subspace and the data manifold, it will be possible to properly represent both the model and the training dataset.
To extract DSVs from the manifold, we leverage the principles of Self-Supervised Learning (SSL). SSL operates under the assumption that a robust model should exhibit augmentation invariance. Inverting this idea, we propose that, when evaluating images within the manifold, an effective model should assess them in an augmentation-invariant fashion. Put differently, the DSVs we identify should remain as DSVs even after augmentation.
5 Experiments
Figure 3: Illustration of entropy change with DSV candidates: We utilized DeepKKT conditions in the selection of DSVs. As training progresses, entropy increases, i.e., the DSVs move toward the decision boundary.
5.1 Implementation Details
We chose the ConvNet architecture (Goodfellow et al., 2016) for pretraining $\Phi(\cdot;\theta)$ on the SVHN dataset (Netzer et al., 2011), a digit dataset with dimensions similar to CIFAR-10 (Krizhevsky, 2009). To synthesize DSVs, we used a large learning rate of 0.1 and a $\rho$ value of 0.0001 were utilized for the SAM optimizer to adhere to the principle of margin maximization (Wang et al., 2023; Foret et al., 2020). For $L_{\text{stat}}$, we set $\alpha_{x}=10^{-3}$, $\alpha_{\lambda}=0.1$, and $\gamma=5$, respectively. For synthesis, it took 15 minutes in a single 1080TI GPU. We leave additional details in Appendix Sec.B
5.2 DSVs: Decoding Black-Box in Deep Learning
DSVs Implicitly Meet Complementary Slackness
As we discussed in Sec. 3.2, the principle of complementary slackness within the KKT conditions suggests that support vectors should be situated on the decision boundary. This aspect was not incorporated explicitly in our model due to computational costs and ambiguity. However, as indicated in Fig. 3, it appears that DSVs implicitly fulfill this condition. Notably, during the training process, we observe an increase in the entropy of DSV candidates.
From a probabilistic perspective, high entropy indicates greater uncertainty. This is a strong indicator that these samples are positioned near the decision boundary, as they are the ones for which the model is least certain about their classification. This aligns with the complementary slackness principle, which posits that support vectors play a crucial role in defining the decision boundary.
Predicting Model Dynamics With DSVs’ Inherent Importance
Figure 4: Correlation between the classwise mean test accuracy versus the sum of $\lambda$’s for training samples belonging to each class. Here, $\rho$ represents the Pearson correlation coefficient.
In SVM, the Lagrangian multiplier $\alpha$, acts as a natural measure to determine which support vector is more important than others. Similarly, DSVs can act a comparable role. Fig. 4 indicates a strong correlation between the sum of $\lambda$’s per each class and actual classwise test accuracy. This phenomenon is intriguing because, despite the model achieving nearly 100% accuracy during training due to overparameterization, DSVs can provide insights into the categorical generalization in the test phase.
Figure 5: Classification result of a mixed image of an original image and a DSV: The model’s prediction follows that of the DSV even though the effect of the DSV is unnoticeable.
DSVs Are Full of Discriminative Information
In Fig. 5, we conducted an experiment by mixing a randomly sampled image from the real dataset with an image from the DSVs. Upon observation, the mixed image is virtually indistinguishable from an image obtained solely from the real dataset. It is noteworthy to highlight this situation resembles that of an adversarial attack (Xu et al., 2019; Moosavi-Dezfooli et al., 2016), yet we did not apply gradient descent to the image; we simply mixed two images. This suggests that the discriminative informational density in a single DSV image is substantially greater than that in a randomly sampled image. The fact that the DSV’s characteristics remained dominant in the classification, underscores the significant role of DSVs in explaining the model’s classification ability.
Figure 6: Model predictions for original versus DSV-informed edited images. Images were altered based on decision criteria derived from DSVs, influencing the model’s prediction.
DSVs Encode the Decision Criteria Visually
Our findings suggest that DSVs for image classification problems not only satisfy the conditions of classical support vectors but also encapsulate a greater volume of visual information. Figure 6 experimentally substantiates our claims and illustrates the practical use of DSVs. Analysis of Figure 1 reveals the decision criteria used for classifying deer, cats, and dogs: 1) DSVs highlight antlers in deer, signifying them as a distinctive characteristic. 2) In cats, pointed triangular ears are a recurring feature in DSVs. 3) For dogs, a trio of facial dots holds significant importance. Using these observations, we altered a deer’s image by erasing its antlers, thereby reducing the model’s confidence in classifying it as a deer, and by reshaping its ears to a pointed contour, which caused the model to misclassify it as a cat. Similarly, by smoothing the ears of a cat image to diminish its classification confidence and then adding antlers or three facial dots, we influenced the model to reclassify the image as a deer or a dog, respectively, according to the introduced features.
DSVs Enable Qualitative Assessment of a Model
This discovery holds significant implications about making models responsible; it introduces a qualitative aspect to assessing model performance. Consider a scenario where we need to differentiate between deer and horses, and the data is collected in summer. The model in our study would be less suitable, as evidenced by Fig. 1, which indicates the model’s reliance on antlers for identifying deer – a feature not present in summer. This shows that DSVs enable us to conduct causal predictions by qualitatively analyzing models, an aspect integral to RAI (Ghallab, 2019). Without DSVs, reliance would be solely on scoring criteria, a potentially hazardous dependence given that model performance is heavily reliant on the quality of data. If the data is biased or contaminated, the model’s ability to make accurate judgments could be compromised. DSVs help mitigate such risks, providing a clearer and more responsible understanding of the model’s decision-making processes.
DSVs: Reconstructing Models With a Minimalist Touch
SVMs utilize support vectors to construct classification models, and this research focuses on DSVs, which serve a similar role to SVM’s support vectors. We have established that DSVs meet the theoretical conditions required by SVM’s support vectors. Based on this, we explored the possibility of reconstructing models using DSVs.
In Table 1, we demonstrate that DSVs can be effectively used to retrain models with a limited amount of training data, aligning with the principles of dataset distillation. Specifically, our experiments were conducted in two approaches, as shown in Fig. 1: coreset selection and image synthesis. For each class, we either generated or selected a single image and these images were used to retrain models from scratch. In terms of image selection, our approach outperformed naive random selection and the forgetting algorithm. This implies that our selection process itself is effective. In the realm of image synthesis, our technique achieved results on par with established dataset distillation methods, surpassing the coreset selection approaches. These findings are noteworthy, especially considering that traditional distillation algorithms depend on strong conditions that they can access to the entire dataset, and have multiple trained models with different initializations. In contrast, our algorithm operates relying solely on a single trained model and does not require any data during the generation process, not even any knowledge about prior dataset as referenced in (Haim et al., 2022). This suggests that the ability of our approach to match the performance of traditional methods, despite not directly utilizing a loss function to enhance the accuracy of retrained models, is a strong indication of the validity of our assertion. It implies that DSVs inherently possess characteristics similar to support vectors in SVMs, as only this could explain their effectiveness in the absence of the direct optimization strategies employed by conventional algorithms.
Table 1: An experimental result of datastet distillation. DD refers to the dataset distillation algorithm in (Wang et al., 2018) and DC refers to the dataset condensation algorithm in (Zhao et al., 2020). In coreset selection, herding refers to  (Chen et al., 2012) and forgetting refers to  (Toneva et al., 2018). DSVs and DSVg represent our method of DSV selection and DSV generation respectively.
coreset selection synthesis
random herding forgetting DSVs DD DC DSVg
14.4 $\pm$ 2.0 21.2 $\pm$ 1.2 13.3 $\pm$ 1.2 17.45 $\pm$ 1.2 27.3 $\pm$ 0.8 28.3 $\pm$ 0.5 25.3 $\pm$ 0.5
DSVs, the First Practical Dataset Distillation
In the context of dataset distillation, DeepKKT can be used as a pioneering algorithm tailored for practical scenarios. Dataset distillation primarily addresses two key issues: 1) Protecting private information through data synthesis, and 2) Reducing the communication load by minimizing the data size. Traditional algorithms, however, face a paradox, particularly when data originates predominantly from edge devices like smartphones (McMahan et al., 2016; Konečný et al., 2016). These algorithms typically require access to the entire dataset, presumably centralized, leading to significant communication overhead and privacy concerns. Contrasting with these approaches, DeepKKT operates independently of any dataset, relying solely on a pre-trained model. This approach eliminates the need for edge devices to store extensive private data. Additionally, due to its lightweight nature, DSVs can be effectively extracted directly on edge devices.
Figure 7: Images of DSVs generated by a model that underwent transfer learning.
DSVs in Practical Settings
[…]
6 Ablation study
Effect of Each Loss
In the synthesis of DSVs, the process starts from noise, which is sampled from a Gaussian distribution. While this initial step resembles the diffusion process in diffusion models (Ho et al., 2020; Song & Ermon, 2019), there is a fundamental difference in the target of the transition. The only difference is its desired distribution. In diffusion models, the transition aims to capture the entire data manifold. However, for DSV synthesis, the focus is specifically on $\partial\mathcal{M}$, the boundary of the data manifold.
A discretized diffusion model with classifier guidance can be described by the following equation:
$x_{t+1}=x_{t}+\epsilon_{t}\cdot\left(\nabla_{x}\log p(x_{t})+\lambda\nabla_{x}\log p(y|x_{t})\right)$ (9)
Our Algorithm 1, which can be formulated as:
$x_{t+1}=x_{t}+\alpha\cdot\left(\nabla_{x}L_{\text{stat}}(A(x))+\beta\nabla_{x}L_{\text{primal}}(A(x)\right)$ (10)
Where $A$ is random augmentation function. In our algorithm, the stationarity condition mirrors the score in diffusion models, while the primal condition aligns with the concept of guidance. However, unlike traditional diffusion models that target the whole data manifold, our process is directed towards navigating from the Gaussian distribution to $\partial\mathcal{M}$, the boundary of the data manifold. This focus on the manifold boundary is a distinctive aspect of our approach, emphasizing the generation of DSVs within the crucial decision boundaries. Fig. 8 also supports our hypothesis that the primal condition alone does not generate meaningful images.. Moreover, while stationarity loss alone can create images with well-structured shapes, it fails to generate accurate samples if the sample labels do not match the targeted label, especially in the absence of primal guidance (classifier guidance). Exploring the intricate relationship between stationarity and diffusion in the context of data manifold navigation presents an exciting avenue for future research.
Figure 8: Results showing DeepKKT images created solely by the primal condition or by the stationary condition.
7 Conclusion
This paper pioneers the definition of support vectors in nonlinear deep learning models, specifically through the introduction of DSVs (Deep Support Vectors). We demonstrated the feasibility of generating DSVs using only pretrained models, without requiring access to the training dataset. To achieve this, we extended the KKT (Karush-Kuhn-Tucker) conditions, to DeepKKT conditions. It can be applied to any deep learning models. The synthesized DSVs effectively encode the decision boundary, offering a visual and intuitive explanation of a model’s decision-making process. Furthermore, we have shown that models can be retrained using DSVs in a manner akin to SVMs.
Societal Impacts
Our paper is closely related to Responsible AI (RAI), particularly in terms of enabling qualitative assessment of models. They provide a visual and intuitive explanation of a model’s decision-making criteria, offering insights into how the model operates that are not only explanatory but also responsible. This approach goes beyond traditional Explainable AI (XAI) methodologies, which typically focus on local explanations for individual inputs and lack the ability to offer a global decision criterion like DSVs. Additionally, as this method is based on model inversion, it ensures safety and privacy. Although the synthesized sets in Fig. 1 might look similar to the selected sets, they do not replicate specific sample features. This is because DSVs represent a more generalized decision boundary, thus avoiding the inclusion of image-specific features. Therefore, DSVs can make all models using logistic loss responsible.
Appendix A intutive explanation of DeepKKT condition
[…]
In DeepKKT, many conditions, except one condition makes sense. For example, primal condition and manifold condition makes sense and dual condition can be regraded as importance sampling. The most counterinutitive part is stationarity section.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{\text{stat}}=D(\theta^{*},-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}))}$ (11)
In this section, we’re to explain the dynamics of overparameterized deep learning, and explain how it is connected to deep learning. Below is an analogy of (Soudry et al., 2018).
A deep learning models follows following ODE:
$w_{t+1}=w_{t}-\alpha\nabla L(x,y;w_{t})$ (12)
Where $\alpha$ is the learning rate and $t$ is the optimization step. $L$ does not go to zero since deep learning models usually exploit a function with a logistic tail, such as the cross-entropy function. Thus, there exists a convergence of the gradient direction $g_{\infty}\coloneqq\hat{\nabla}L$. There also exists a time $T$ where the gradient direction converges to $g_{\infty}-\varepsilon$ for sufficiently small $\varepsilon$. As illustrated in Fig. 9, $w$ moves toward the direction of $-g^{\infty}$. Therefore, $\hat{w}_{\infty}\approx-g_{\infty}$.
This is what stationarity condition wants to seek for. The direction of $g_{\infty}$, by using only few support vectors.
Appendix B Full Implementation Details
For $L_{\text{stat}}$, we set $\alpha_{x}=10^{-3}$, $\alpha_{\lambda}=0.1$, and $\gamma=5$, respectively. This was due to the differing scales. The actual computed loss for stationarity was ten times higher than for primal loss. To satisfy the manifold condition in reconstructing DSVs, we implemented techniques from the diffusion model (Rombach et al., 2021). This involved synthesizing data at a low resolution, then upscaling it to a higher resolution. We initially downscaled the image twice, then upscaled it, continuing with optimization. To reduce the impact of outliers, we choose the $L_{1}$ metric and apply the square root to the gradient term when measuring distance.
To obtain the pretrained weight $\theta^{*}$, we chose the ConvNet architecture (Goodfellow et al., 2016), a common choice in deep learning. This architecture includes sequential convolutional layers followed by max pooling, and a single fully-connected layer for classification. The learning rate was set to $10^{-3}$ with a weight decay of $0.005$ using the Adam optimizer. Additionally, we employed flipping and cropping techniques, with settings differing from those used for DSVs reconstruction to ensure fair comparison. For pretraining $\Phi$ on the Street View House Numbers (SVHN) dataset (Netzer et al., 2011), a digit dataset with dimensions similar to CIFAR-10 (Krizhevsky, 2009), we exclusively trained the fully-connected layer of the CIFAR-10 pre-trained ConvNet. This approach resulted in a training accuracy of 80%. When retraining the model with DSVs, we used large learning rate of 0.1 and a $\rho$ value of 0.0001 were utilized for the SAM optimizer to adhere to the principle of margin maximization (Wang et al., 2023; Foret et al., 2020). For DSVs synthesis, we used a single 1080TI GPU and it took 15 minutes.
Figure 10: Images of DSV candidates (Selection).
Figure 11: Images of various DSVs (Synthesized) – Cat, Ship, Dog, Car, Truck, Bird, Horse, Deer, Plane, Frog.
Figure 12: Synthesized images on other architecture.